OpenStudy (anonymous):

MEDAL FOR ALL AND ANY ANSWERS!!!!!!!!!!!!!!!! 4x + 3y 36..... give me an ordered pair that's all i need.

OpenStudy (anonymous):

is that supposed to be +36?

OpenStudy (anonymous):

Is tht 4x+3y=36?

OpenStudy (anonymous):

y = -4/3x + 12

OpenStudy (anonymous):

its supposed to be 4x+ 3y \[\ge\] 36

OpenStudy (anonymous):

(5, 6)

OpenStudy (anonymous):

It doesn't have a solution...

OpenStudy (anonymous):

It has a lot of solutions. 4 times x + 3 times 7 has to be greater or equal to 36. (1000, 360) is a solution.

OpenStudy (anonymous):

i mean 3 time y. sorry

OpenStudy (anonymous):

Well when i do the problem i'm not getting tht and where did u get 7?

OpenStudy (anonymous):

*oh okay

OpenStudy (anonymous):

4 times 5 (x) is 20 + 3 times 6 (y) is 18 so 38 is greater or equal to 36 so it works.

OpenStudy (anonymous):

oh okay i get it now

OpenStudy (anonymous):

cool, yeah i thought it was a different problem at first too cause it looked a bit weird xD thanks.

OpenStudy (anonymous):

yea same here thts y i didn't get a soution then u just explained it to me and i gert it now...nd ur welcum

OpenStudy (anonymous):

*get